# A Method for Localization of Cellular Users from Call Detail Records

### v1

@@ I Introduction
The principal contribution of this paper is a method for CDR-based localization in which the RPL shape is derived from a model of the antenna pattern via the Friis Transmision Equation, using a well-established model for terrestrial radio path loss that is applicable to outdoor macrocells at UHF frequencies. This shape is then scaled to account for BS spatial density, which bypasses the problem of estimating sector hand off boundaries. The efficacy of the method is demonstrated using drive test measurements.
@@ II RPL Estimation Method
It is not normally possible to calculate $r_{max}$ in an CDR-based localization problem. Among the many difficulties is that $P_{T}$ is typically unknown and varies dynamically due to power control, $G_{R}/P_{R,min}$ varies over orders of magnitude depending on the phone and user behaviors, and the propagation parameters $n$ and $r_{b}$ are both unknown and potentially vary with azimuth $\phi$.

### v2

@@ I Introduction
The principal contribution of this paper is a method for CDR-based localization in which the RPL shape is derived from a model of the antenna pattern via the Friis Transmision Equation, using a well-established model for terrestrial radio path loss that is applicable to outdoor macrocells at UHF frequencies. This shape is then scaled to account for BS spatial density, which bypasses the problem of estimating sector hand off boundaries. To the best of our knowledge, this is a novel method for RPL estimation. The efficacy of the method is demonstrated using drive test measurements. While insufficient information exists for a meaningful quantitative comparison of performance to other methods addressed in the previous paragraph, the method and results reported here are suitable for future “apples-to-apples” comparisons to both (1) the proposed method employed in other scenarios, and (2) other methods when results are reported in sufficient quantitative detail.
@@ II RPL Estimation Method
It is not normally possible to calculate $r_{max}$ in an CDR-based localization problem. Among the many difficulties is that $P_{T}$ is typically unknown and varies dynamically due to power control, the received power may differ from the mean power due to local shadowing and diffraction, $G_{R}/P_{R,min}$ varies over orders of magnitude depending on the phone and user behaviors, and the propagation parameters $n$ and $r_{b}$ are both unknown and potentially vary with azimuth $\phi$.
@@ V Conclusions
A limitation of the method is that the proposed RPL shape cannot account for persistent deviations from the mean power density due to shadowing and diffraction associated with buildings and terrain. This may explain some fraction of the errors observed in the field experiment reported here, and so this would be an appropriate topic for future work.